Find the mass and center of gravity of the cylindrical solid that has density δ(x, y, z) = h − z and is enclosed by

x

2 + y

2 = a

2

, z = 0, and z = h


1
Expert's answer
2021-11-30T11:20:21-0500

"mass,m=\int\int\int\delta(x,y,z)dV\\
=\int_{x=-a}^a\int_{y=-a}^a\int_0^h(h-z)dzdydx\\
=\int_{x=-a}^a\int_{y=-a}^a[hz-\frac{z^2}{2}]_0^hdydx\\
=\frac{h^2}{2}\int_{x=-a}^a\int_{x=-a}^a dydx\\
=\frac{h^2}{2}\int_{x=-a}^a[y]_{-a}^a dx\\
=a{h^2}\int_{x=-a}^a dx\\
=a{h^2}[x]_{-a}^a\\
=2a^2h^2\\

\text{Now, calculate moments}\newline

M_{xy}=\int\int \int z\delta dV
=\int_{x=-a}^a\int_{y=-a}^a\int_0^h hz-z^2dzdxdy\\
=\int_{x=-a}^a\int_{y=-a}^a[\frac{hz^2}{2}-\frac{z^3}{3}]_0^hdydx\\
=\frac{h^3}{6}\int_{x=-a}^a\int_{x=-a}^a dydx\\
=\int_{x=-a}^a[y]_{-a}^a dx\\
=a\frac{h^3}{6}\int_{x=-a}^a dx\\
=a\frac{h^3}{6}[x]_{-a}^a\\
=2a^2\frac{h^3}{6}
\\

M_{yz}=\int\int \int x\delta dV
=\int_{x=-a}^a\int_{y=-a}^a\int_0^h x(h-z)dzdxdy\\
=\int_{x=-a}^a\int_{y=-a}^ax[hz-\frac{z^2}{2}]_0^hdydx\\
=\frac{h^2}{2}\int_{x=-a}^a\int_{x=-a}^a xdydx\\
=\frac{h^2}{2}\int_{x=-a}^ax[y]_{-a}^a dx\\
=a{h^2}\int_{x=-a}^a xdx\\
=a{h^2}[\frac{x^2}{2}]_{-a}^a\\
=0\\


M_{xz}=\int\int \int y\delta dV
=\int_{x=-a}^a\int_{y=-a}^a\int_0^h y(h-z)dzdxdy\\
=\int_{x=-a}^a\int_{y=-a}^ay[hz-\frac{z^2}{2}]_0^hdydx\\
=\frac{h^2}{2}\int_{x=-a}^a\int_{x=-a}^a ydydx\\
=\frac{h^2}{2}\int_{x=-a}^a[\frac{y^2}{2}]_{-a}^a dx\\
=a{h^2}\int_{x=-a}^a 0dx\\
=0\\

\text{Center of gravity is given by}(x,y,z)then,\\
x=\frac{M_{yz}}{m}=\frac{0}{2a^2h^2}=0,\\
y=\frac{M_{xz}}{m}=\frac{0}{2a^2h^2}=0,\\
z=\frac{M_{xy}}{m}=\frac{2a^2\frac{h^3}{6}
}{2a^2h^2}=\frac{h}{6}
\\"